Calculate the heat released when 68.0 gram of steam at 1240C is converted to water

    at 450C


1
Expert's answer
2022-05-18T09:17:03-0400

released can be calculated by the following formula.

"q=ms\Delta t"

m = Mass of sample

s = Specific heat of steam

"\Delta t" =Temperature change

From the given,

Mass of sample = 68.0 g

Specific heat of steam = 1.99 J/(g°C)

"\Delta t=124-100=24°C"

"q=68x1.99x24=3247 J"

s1 - Specific heat of water, 4.184 J

"\Delta t=100-45=55°C"

"q_1=68x4.184x55=" 15648

Q=q+q1=3247+15648=18895 J